OpenStudy (anonymous):

A cereal box is 8 inches long, 3 inches wide, and 12 inches tall. (a) Volume? (b) Surface area?

OpenStudy (anonymous):

Total Surface Area = 2(Areahw) + 2(Areahl) + 2(Areawl) Volume = (h)(w)(l)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

where h is height w is wide and l is length

OpenStudy (anonymous):

did u get it?;

OpenStudy (anonymous):

not really im tierds i cant think

OpenStudy (anonymous):

The total surface area is made up of three pairs of sides for a total of six sides. We must find Areahw, which is the area of the side that is h by w. Areahw = (h)(w) We must find Areahl, which is the area of the side that is h by l. Areahl = (h)(l) We must find Areawl, which is the area of the side that is w by l. Areawl = (w)(l) Each of the above areas are present on two sides of the box. So, the total surface area is sum of twice each of these areas, that is: Total Surface Area = 2(Areahw) + 2(Areahl) + 2(Areawl

OpenStudy (anonymous):

gtg now felt asleep

OpenStudy (anonymous):

ill fan and medal i just need answer

OpenStudy (anonymous):

u can put the value of h w l and u would get the correct answer...i really have to go bye///ttyl tc